Suppose that in 2011, potential GDP for BloomJack Island is $85,000, real GDP is $75,000, and potential GDP grows at a rate of 2% per year.

  1. If real GDP is $78,000 in 2012, using Okun's law, calculate the cyclical rate of unemployment.

b. If real GDP is $83,000 in 2013, using Okun's law, calculate the cyclical rate of unemployment.
Answer #1

According to the law, we have output gap = -2*unemployment gap implying that when unemployment rate exceeds its natural rate by 1 percent (there is a cyclical unemployment of 1 percent) GDP grows 2 percent less than its potential.

a) In 2012, potential GDP is 85000(1 + 2%) and real GDP is 78000.

This implies the output gap is (78000 - 85000(1 + 2%))*100/(85000(1 + 2%)) = -10.03%.

This gives cyclical unemployment = 10.03/2 = 5.02%

b) In 2013, potential GDP is 85000(1 + 2%)^2 and real GDP is 83000.

This implies the output gap is (83000 - 85000(1 + 2%)^2)*100/(85000(1 + 2%)^2) = -6.14%.

This gives cyclical unemployment = 6.14/2 = 3.07%